OpenStudy (anonymous):

find the derivative y=1-ln(1/x)

OpenStudy (anonymous):

use the chain rule f(x)=-ln(x) g(x)=1/x d/dx f(g(x)) = f'(g)*g'(x)

OpenStudy (ybarrap):

$$ \cfrac{d}{dx}(1)=? $$ That should be easy $$ \ln x=\int_1^x\cfrac{1}{x}dx $$ So the derivative of the integral is what? Put these together. Problem solved.